Your company is considering purchasing a new machine. It will take three years to put into operation, but at the end of the third year it will increase company income by $10,000, increasing by $1,000 per year for the following five years. At the end of the eighth year it will be worn out and have no scrap value. If the cost of capital to the company is 15%, what is the most it is worth paying for the machine now? (5 pts)
Answer #1

ANSWER:

Increase in company's income = $10,000

further increase per year for following 5 years = $1,000

i = 15%

pw = income in 3rd year(p/f,i,n) + income in 4th year(p/f,i,n) + income in 5th year(p/f,i,n) + income in 6th year(p/f,i,n) + income in 7th year(p/f,i,n) + income in 8th year(p/f,i,n)

pw = 10,000(p/f,15%,3) + 11,000(p/f,15%,4) + 12,000(p/f,15%,5) + 13,000(p/f,15%,6) + 14,000(p/f,15%,7) + 15,000(p/f,15%,8)

pw = 10,000 * 0.6575 + 11,000 * 0.5718 + 12,000 * 0.4972 + 13,000 * 0.4323 + 14,000 * 0.3759 + 15,000 * 0.3269

pw = 6,575 + 6,289.8 + 5,966.4 + 5,619.9 + 5,262.6 + 4,903.5

pw = 34,617.2

The most that the company should pay for the machine is $34,617.2